OpenStudy (anonymous):

Find the solution set to each equation. x+7/x+4 = x+1/x-2 Please help!!

OpenStudy (anonymous):

There are no real solutions for x.

OpenStudy (anonymous):

Oh this equation is no solution? cause it'd be x+7=0/x+4=0 right

OpenStudy (anonymous):

Yes. well, which class are u and where are u studying?

OpenStudy (anonymous):

West Coast University. It's called Algebra for college students. I have this huge test at 7am :(

OpenStudy (anonymous):

All the very best.

OpenStudy (anonymous):

I have another question, have time? & thanks!

OpenStudy (anonymous):

Please ask. If I acn, I will help u.

OpenStudy (anonymous):

It's really easy but I keep messing up on one part, & just need someone to explain it. Reduce to the lowest terms. (2-x)(x+1)(x+3) I know for the (x+1)(x+3) it is x^2+3x+1x+3

OpenStudy (anonymous):

After that I just get stuck and mess up my #'s so do I distribute the (2-x) to the original or to x^2+3x+1x+3

OpenStudy (anonymous):

Answer= -x^3 - 2x^2 + 5x - 6 (After u get x^2+4x+3, multiply this by 2-x)

OpenStudy (anonymous):

\[(2-x)(x+1)(x+3)=(2-x)(x^2+4x+3)=2x^2+8x+6-x^3-4x^2-3x=-x^3-2x^2+5x+6\]

OpenStudy (anonymous):

oh okay so my question do I start off with 2 for example 2(x^2)+2(4x)+2(5x)-2(6) get what I'm saying, that part totally throws me off, I know it's easy but I'm stuck :/

OpenStudy (anonymous):

I got it!